WHat is the IQR of the set 5,5,6,6,6,7, 7, 7, 7, and 8 ?
Answer #1

here the first half data set : 5,5,6,6,6

second half data set : 7,7,7,7,8

Q2 = median of data set = (6+7)/2 = 6.5

from 1st data set: Q1 = 6 = median of 1st half of  data set

from 2nd data set : Q3 = 7 = median of 2nd half of data set

then, IQR = Q3 - Q1 =7 - 6 =1